# Application of the microscopic optical potential of chiral effective field theory in astrophysical neutron-capture reactions

### v1

@@ 4 Conclusions
In this work, we apply the microscopic global optical potential of chiral effective field theory (the WLH potential) to the study of astrophysical neutron-capture reaction for the first time. By comparing the neutron-capture cross sections calculated using the WLH potential with experimental data for the nuclei ${}^{61}\text{Ni}$, ${}^{122}\text{Sn}$, and ${}^{176}\text{Lu}$, we find that the theoretical predictions show good agreement with the experimental results. Furthermore, for all experimentally known MACS with $26\leq Z\leq 83$, the WLH potential also provides a satisfactory description. We employ the $f_{rms}$ deviation to assess the predictive accuracy of the WLH potential for each sample of the parameter set, and find that the associated uncertainty lies within a reasonable range. These results demonstrate the capability of the WLH potential to reliably reproduce experimental data and establish a foundation for its application to unstable neutron-rich nuclei.
For neutron-rich nuclei, to efficiently propagate parameter uncertainties to reaction rates and reduce computational cost, we adopt the BFMC method, using the $f_{rms}$ deviation to constrain parameter sampling. With this approach, we calculate the neutron-capture rates for nuclei ${}^{122}\text{Sn}$, ${}^{132}\text{Sn}$, and ${}^{150}\text{Sn}$ at different temperatures, as well as the rates for Sn isotopes as a function of neutron number at a temperature of $T=1$ GK. It is found that the $f_{rms}$ constraint effectively controls the propagation of uncertainty, and a sharp increase in uncertainty is observed around $N=87$. A further comprehensive calculation, covering all nuclei with $26\leq Z\leq 84$ between the valley of stability and the neutron drip line at $T=1$ GK and constrained by $f_{rms}\leq 1.56$, shows that the WLH potential provides a good description with relatively small uncertainties for nuclei with isospin asymmetry $\delta<0.28$. However, for nuclei with $\delta>0.28$, the uncertainty in the reaction rates increases significantly. This highlights the critical role of isospin dependence in optical potentials and suggests that future developments of the WLH potential may pay greater attention to this aspect.
@@ 5 Acknowledgements
This work is supported by the National Key Research and Development program (MOST 2022YFA1602304), the National Natural Science Foundation of China (Grants No. 12335009, 12175156, and 12375122), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB34020200), and the Fundamental Research Funds for the Central Universities (Grant No. B240201048). Y. X. acknowledges support from the Romanian Ministry of Research, Innovation and Digitalization under Contract No. PN 23 21 01 06 and the ELI-RO project with Contract ELI-RORDI-2024-008 (AMAP), and support from the IAEA CRP on ”Updating and Improving Nuclear Level Densities for Applications (Project No. F41034)”.

### v2

@@ 4 Conclusions
In this work, we apply the microscopic global optical potential of chiral effective field theory (the WLH potential) to the study of astrophysical neutron-capture reaction for the first time. By comparing the neutron-capture cross sections calculated using the WLH potentials with experimental data for the nuclei $^{61}\text{Ni}$, $^{122}\text{Sn}$, and $^{176}\text{Lu}$, we find that the theoretical predictions show good agreement with the experimental results. Furthermore, for all experimentally known MACS with $26\leq Z\leq 83$, the WLH potentials also provide a satisfactory description. We employ the $f_{rms}$ deviation to assess the predictive accuracy of the WLH potential for each sample of the parameter set, and find that the associated uncertainty lies within a reasonable range. These results demonstrate the capability of the WLH potentials to reliably reproduce experimental data and establish a foundation for its application to unstable neutron-rich nuclei.
For neutron-rich nuclei, to efficiently propagate parameter uncertainties to reaction rates and reduce computational cost, we adopt the BFMC method, using the $f_{rms}$ deviation to constrain parameter sampling. With this approach, we calculate the neutron-capture rates for nuclei $^{122}\text{Sn}$, $^{132}\text{Sn}$, and $^{150}\text{Sn}$ at different temperatures, as well as the rates for Sn isotopes as a function of neutron number at a temperature of $T=1$ GK. It is found that the $f_{rms}$ constraint effectively controls the propagation of uncertainty, and a sharp increase in uncertainty is observed around $N=87$. A further comprehensive calculation, covering all nuclei with $26\leq Z\leq 84$ between the valley of stability and the neutron drip line at $T=1$ GK and constrained by $f_{rms}\leq 1.56$, shows that the WLH potentials provide a good description with relatively small uncertainties for nuclei with isospin asymmetry $\delta<0.28$. However, for nuclei with $\delta>0.28$, the uncertainty in the reaction rates increases significantly. This highlights the critical role of isospin dependence in optical potentials and suggests that future developments of the WLH potential may pay greater attention to this aspect.
The MACS of neutron capture is directly related to the astrophysical cross section, and a large amount of experimental data can be utilized. Other experimental observables, such as elastic scattering data and the neutron strength function, may be more sensitive to the neutron optical potentials and thus offer stronger constraints to the parameter sampling. Following the sensitivity check of neutron optical potentials to the neutron strength function presented in this work, a systematic quantitative study about how the neutron optical potentials can be constrained by different experimental observables would be a worthwhile direction in future.
Meanwhile, in the WLH potential, nuclear deformation is neglected, and its effects are not included in the present study. However, since nuclear deformation plays an important role in reproducing nucleon–nucleus scattering experimental data, incorporating this effect into the WLH potential to describe neutron-capture reactions would be worthwhile in future work.
@@ 5 Acknowledgements
This work is supported by the National Key Research and Development program (MOST 2022YFA1602304), the National Natural Science Foundation of China (Grants No. 12335009, 12175156, and 12375122), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDB34020200), and the Fundamental Research Funds for the Central Universities (Grant No. B240201048). Y. X. acknowledges support from the Romanian Ministry of Research, Innovation and Digitization, CNCS-UEFIS-CDI, project number PN-IV-P1-PCE-2023-0384 within PNCDI IV, the Romanian Ministry of Research, Innovation and Digitalization under Contract No. PN 23 21 01 06, the ELI-RO project with Contract ELI-RORDI-2024-008 (AMAP), the IAEA Coordinated Research Project on Updating Nuclear Level Densities for Applications (F41034) under Contract No. 28638.